# Legally Binding but Unfair? Towards Assessing Fairness of Privacy Policies

### v1

@@ III-B Dimensions of Fairness
Ethics and Morality: Art. 5(1) point (a), 6(2)-(3) GDPR and Recital 71 states fairness as a fundamental principle, grounded in ethics and morality. The GDPR strives to avoid potential harm to data subjects through safeguards, and represents procedural fairness as a central element of its fairness conception [67]. This means users should be protected from vulnerabilities and risk. Interests between the data subject and service provider should be fairly balanced [67].
@@
We are open to considering more dimensions of fairness in the future. In the following sections, we explore our three fairness dimensions. We propose approaches to assess indicators for each dimension, we show the results of preliminary experiments, and we discuss how this brings forward fairness in privacy policies.
IV Informational Fairness
@@
Our assessment of informational fairness borrows from the approaches described in Section II-B1.
IV-A Approach
@@
Informational fairness in privacy policies is related to the completeness of a privacy policy and the comprehensibility and readability of the vocabulary, sentences, and document structure for all social or demographic groups.
Completeness
@@
To get an estimate of the completeness of privacy policies, we propose to use supervised learning, as described in [61]. We want to identify the relevant 56 metadata types and then assess the embedding space to judge completeness.
Fairness on word level
@@
Anglicisms, complex words, tech jargon, etc. induce a socioeconomic, nationality, ableist, and age-related bias, causing unfairness due to discrimination. We propose lexical filtering, either with an English language dictionary, a word-to-word translation library, or a customized dictionary, to detect tech jargon and anglicisms. A frequency-based dictionary [68, 69] allows us to estimate the proportion of words in a policy, that are not in common usage with a threshold-based approach. We can estimate the use of complex or ambiguous words, when we use a translation service to translate words into another language and back, and measure the proportion of words that remain unchanged. However, the translation service might be affected by biases [70] and linguistic limitations, e.g., when translating a gendered language to a gender-neutral one and back. We also propose the Bradley-Terry statistical model with embeddings to assess lexical ambiguity [36].
Fairness on sentence level
@@
Incomprehensible policies lead to discrimination and unfairness. We propose to estimate the readability of the sentences using statistical text metrics and existing linguistic measures. A straightforward text metric is the number of words in the policy. A higher number indicates less readable policies. The Flesch Reading Ease (FRE) [37] and the TAASSC 2.0 [40] are well-recognized linguistic metrics for readability. The coherence of sentences can be assessed with the DiscoScore metric [41].
Inclusive document structure
@@
An inclusive policy structure mitigates socioeconomic, nationality, and ableist bias and improves fairness. Again, we propose statistical and linguistic measures. An inclusive policy is indicated by short section headings or lengths per paragraph, or a shallow split of the policy into sections (structural depth). To assess how well the headings semantically fit with the whole section text of a policy, embeddings can be used [71, 72].
IV-B Preliminary Results
@@
As a test case, we use the German Top-100 most visited web shops. Our data set contains the canonical form of 618 German privacy policies over eight years, starting in 2016. For acquisition, preprocessing, cleansing, and a description of the data set, see [58]. We conducted all experiments in German, as this is the language of the data set, but translated the results to English for this paper.
Concerning fairness on word level, we exemplarily selected eight policies, one for each year in the data set, and conducted lexical filtering. A simple dictionary-based approach to filtering for potential anglicisms produces many false positives. Many German stop words in NLTK [73] overlap with English words from SpaCy’s [74] English dictionary and can be removed. We then used a word-to-word translator [75] with a smaller vocabulary than SpaCy or NLTK, to narrow down our list of anglicisms, and measured the length of this list. In our eight policies, we found 26 anglicisms on average, which impairs informational fairness by discriminating against demographics, where anglicisms are not commonly used.
@@
TABLE I: Overview of surface measures
@@
words / policy
4809.59
paragraphs / policy
149.63
words / paragraph
32.14
headings / policy
33.82
6.15% of policies no heading
different type headings
2.85
words / heading
4.83
lists
91.59 % of policies
other text formatting
strong: 71.52 % of policies
italics: 15.37 % of policies
Table I provides statistics on the document structure of all 618 policies. The average policy is very long. A dyslexic, who needs $\approx$2 minutes for 250 words [76, 77], would need to spend almost 165 minutes reading the longest policy. The average reading time per policy for an average reader [76] would be just under 20 minutes.
@@
IV-C Discussion
Text statistics, FRE and, to some extent, formatting have already been measured for privacy policies aiming at transparency [58]. However, existing work did not use this information to assess key aspects of fairness. While informational fairness has previously been assessed for the predictions of machine learning models [24], privacy policies are a different problem. Privacy policies are intended to balance the information asymmetry between provider and user (completeness), and add the aspect of potential discrimination due to a lack of inclusiveness in the presentation of information.
V Representational Fairness
Our approach regarding representational fairness uses the methods outlined in Section II-B2 to assess the ability-related, socioeconomic, age-related, political, nationality-related, gender-related, sexuality-related, and cultural social / demographic axes.
V-A Approach
Assessing representational fairness requires descriptor terms that capture various demographics (cf. Sec. II). Based on the descriptor terms, a quantitative and semantic assessment can be carried out.
Fairness descriptor terms
We propose to carefully select existing descriptor terms, e.g., for gender [78], or via translation service from other languages [44]. Finally, we propose to use a multilingual LLM to filter the resulting set, generate contextualized translations, or automatically generate policy-specific descriptor terms.
Quantitative and semantic fairness assessment
The descriptor terms enable us to estimate how often different groups are represented in a policy. Semantically, we propose to use the descriptor terms to assess the invariance of sentiment toward changed demographic group membership by replacing existing entities with changed ones of varying demographic groups. Finally, we propose to leverage the SAME metric [46] for bias detection. We want to adapt approaches for toxicity detection [47], and debiasing the detection itself [48].
V-B Preliminary Results
Again, our test case is based on German language privacy policies from the Top-100 German web shops. With this series of experiments, we evaluated the entire data set of 618 policies. We translated the descriptor terms from HolisticBias [44] to German with deep-translator [79] utilizing the Google Translator API, and we manually removed homonyms and other artifacts from the translation. We appended terms from the categories “nationality” and “political” that are specific for a German context, and we integrated German gender descriptor terms [78]. We perform a quantitative assessment.
@@ V-B Preliminary Results
V-C Discussion
To the best of our knowledge, representational fairness has not yet been assessed for privacy policies. We have only investigated quantitatively as a first step. For more nuanced and accurate findings we want to assess semantics as a next step. This addresses false positives due to homonyms and has been used effectively in recent research (cf. Sec.  II-B2).
VI Ethics and Morality
This assessment uses approaches from Section II-C to screen the processes, rights, and obligations declared in a privacy policy.
VI-A Approach
@@
We propose to train a specific vulnerability classifier, similar to [63]. Based on that, the proportionality of a policy depends on the business activities of the data holder. The assessment of proportionality requires weighing the difficult-to-quantify reasons for a case-specific process against its induced vulnerabilities. This raises challenges for typical ML classification models, so we suspect an ethicist’s judgment to be superior (cf. Section II). That’s why we propose utilizing an LLM as a new approach to be tested. We use the classification result of the vulnerability classifier to prompt the LLM.
General Ethics
@@
We also suggest testing general ethics issues with an LLM, using a broad framing of the prompt. Because this is an explorative approach, we propose to request the LLM to state the criteria of its assessment first, and to quantify them on a five-point Likert scale in a subsequent step. We know from prompt engineering [57] that this can be achieved by using the entire policy as context and a task description in the prompt, and by asking the LLM to reflect its result. As a supplement, we suggest approaches for sentiment analysis [80] on the LLM outputs, to get an approximate overall proxy for the ethics of a policy.
VI-B Preliminary Results
@@
Again, we use the German language policies from the German Top-100 web shops. For the time being, we focus on general morality and leave aside the other aspects. We created a prompt with a broad scope (see Appendix A), and tested it with GPT4turbo [53] on eight policies, one for each year in the data set. We asked the LLM not only to make a qualitative assessment, but also to rate it on a five-point Likert scale. Due to the lack of space, we only report the Likert ratings (Table II).
TABLE II: Overview of LLM ethics assessments
Criteria
Runs
Policies
Average
Score
transparency (precise, complete, comprehensible)
38
8
3.20
data subject control and autonomy (user rights, freedom of choice)
39
8
3.35
data minimalization and purpose binding
26
8
3.06
data usage (surveillance concerns, automated decisions)
21
8
2.71
data storage and deletion
8
4
3.38
data protection and security (avoiding misuse or leakage)
30
8
3.38
(international) data
transmission and sharing (third-parties, where is what transferred)
27
8
2.74
compliance with data protection regulation (GDPR, relevant countries data protection regulation and standards)
9
7
3.56
communication and changes (timeliness, changes, risks)
8
5
2.63
protection of minors
4
2
3.50
fairness (no discrimination, fair use)
7
4
3.14
assurance, accountability and governance (compliance, reliability, assurance, ability to react)
9
6
3.22
@@
As the first column of Table II shows, we can aggregate the LLM assessments into 12 distinct criteria for the ethics of privacy policies. As desired, the criteria capture general ethical requirements on a broad scope and go beyond the GDPR, e.g., regarding the protection of minors. An LLM answers each prompt with a new internal seed, comparable to an expert asked to make an assessment from a different point of view. For this reason, we let the LLM evaluate each prompt in five runs. The second and third columns of Table II show in how many runs and policies the criteria appear. Some criteria just appeared in one of the runs for a policy. While integrating a fixed set of criteria into the prompt would enable a standardized assessment, running the LLM multiple times results in different aspects being considered, which has some merit for exploration. The fourth column shows the Likert-scores between 1 (worst) and 5 (best). Our most alarming findings are the usage of data, the (international) transmission, and sharing of data, as well as communication and changes.
VI-C Discussion
@@
Existing work has not automated ethics assessment so far, i.e., an LLM for ethics evaluation of privacy policies is a novel approach. Our findings are in line with related work (cf. Sec.  II-D) regarding problematic data use. The GDPR’s principles are the legal backbone for ethics in privacy policies. We found these principles represented in the ethics criteria, which we identified automatically with GPT4turbo. As we exemplarily investigated eight policies, our list of criteria is just a starting point. There may be further criteria worth considering.
VII Use Cases and Applications
@@
Our approach enables researchers to automatically scan a large corpus of privacy policies for such issues. This allows to obtain an overview of how widespread which kind of fairness issue is in privacy policies. Due to the ties between fairness and the GDPR, this might also help authorities to single out privacy policies that are problematic from a legal point of view. Our data set allows to analyze how such issues have developed or subsided over time. This provides an opportunity for researchers to investigate whether changes in legislation or major privacy events have led to fairer or less fair policies.
Writing a fair policy is challenging. In the case of AI applications, for example, it is difficult even for experts to assess the personal risks associated with the use of personal data. Our approach allows to automatically identify aspects, that are associated with various kinds of bias, discrimination, etc. It might be useful to integrate this either into existing generators for privacy policies or to develop a stand-alone tool, to support data protection officers in writing complex, but fair privacy policies. Also, our work can be used as a foundation for data assessment for legal NLP models trained on privacy policies.
VIII Conclusion
@@
Appendix A LLM Prompt for ethics
This prompt template for an LLM assesses general morality with a broad scope. The prompt template consists of a system prompt (Figure 1) that defines the general behavior of the LLM, and a prompt (Figure 2) that contains the actual input data. Square brackets are replaced with the privacy policy. For better understanding, we have translated our German prompt template to English.
You are a critically scrutinizing, experienced expert in an ethics council on issues around fairness and justice and are committed to consumer protection for fairness. You have 20 years of experience in ethical issues around consumer protection, data protection, cybersecurity and related fairness aspects. You have already provided precise expert opinions in many court proceedings. Your task is to uncover aspects in privacy policies that are ethically questionable. You must clearly explain your thoughts and how you came to your conclusion. Explain step by step! Also disclose the ethical and moral principles on which you base your judgment. In particular, point out potential for improvement or risks to users. IMPORTANT: Also rate the privacy policy for each of your criteria on a 5-point Likert scale. Check whether the following privacy policy is fair and ethical towards its users. The user wants a critical review and an assessment of vulnerabilities in privacy policies.
Figure 1: System Prompt
The privacy policy: [Privacy Policy]
Figure 2: Prompt

### v2

@@ 3.2. Dimensions of Fairness
Ethics and Morality: Art. 5(1) point (a), 6(2)-(3) GDPR and Recital 71 states fairness as a fundamental principle, grounded in ethics. The GDPR strives to avoid potential harm to data subjects through safeguards, and represents procedural fairness as a central element of its fairness conception (Malgieri, 2020). This means users should be protected from vulnerabilities and risk. Interests between the data subject and service provider should be fairly balanced (Clifford and Ausloos, 2018; Malgieri, 2020).
@@
We are open to considering more dimensions of fairness in the future. We want to highlight that our three fairness dimensions are just motivated by the legal foundation and go beyond legal compliance. We propose approaches to assess indicators for each dimension, we show the results of preliminary experiments, and we discuss how this brings forward fairness in privacy policies.
4. Informational Fairness
@@
Our assessment of informational fairness borrows from the approaches described in Section 2.2.1.
4.1. Approach
@@
Informational fairness in privacy policies is related to completeness of a privacy policy and comprehensibility and readability of vocabulary, sentences, and document structure for protected groups.
@@
Completeness: To get an estimate of the completeness of privacy policies, we propose to use supervised learning, as described in (Amaral et al., 2021). We want to identify the relevant 56 metadata types and then assess the embedding space to judge completeness.
@@
Fairness on word level: Anglicisms, complex words, tech jargon, etc. induce a socioeconomic, nationality, ableist, and age-related bias, causing unfairness due to discrimination. We propose lexical filtering, either with an English language dictionary, a word-to-word translation library, or a customized dictionary, to detect tech jargon and anglicisms. A frequency-based dictionary (Schäfer, 2015; Schäfer and Bildhauer, 2012) allows us to estimate the proportion of words in a policy, that are not in common usage with a threshold-based approach. We can estimate the use of complex or ambiguous words, when we use a translation service to translate words into another language and back, and measure the proportion of words that remain unchanged. However, the translation service might be affected by biases (Savoldi et al., 2021) and linguistic limitations, e.g., when translating a gendered language to a gender-neutral one and back. We also propose the Bradley-Terry statistical model with embeddings to assess lexical ambiguity (Liu et al., 2022).
@@
Fairness on sentence level: Incomprehensible policies lead to discrimination and unfairness. We propose to estimate the readability of the sentences using statistical text metrics and existing linguistic measures. A straightforward text metric is the number of words in the policy. A higher number indicates less readable policies. The Flesch Reading Ease (FRE) (Flesch, 1948) and the TAASSC 2.0 (Kyle et al., 2021) are well-recognized linguistic metrics for readability. The coherence of sentences can be assessed with the DiscoScore metric (Zhao et al., 2023).
@@
Inclusive document structure: An inclusive policy structure mitigates socioeconomic, nationality, and ableist bias and improves fairness. Again, we propose statistical and linguistic measures. An inclusive policy is indicated by short section headings or lengths per paragraph, or a shallow split of the policy into sections (structural depth). To assess how well the headings semantically fit with the whole section text of a policy, embeddings can be used (Kenter and De Rijke, 2015; Han et al., 2021).
4.2. Preliminary Results
@@
As a test case, we use the German Top-100 most visited web shops. Our data set contains the canonical form of 618 German privacy policies over eight years, starting in 2016. For acquisition, preprocessing, cleansing, and a description of the data set, see (Bartelt and Buchmann, 2024). We conducted all experiments in German, as this is the language of the data set, but translated the results to English for this paper.
Concerning fairness on word level, we exemplarily selected eight policies, one for each year in the data set, and conducted lexical filtering. A simple dictionary-based approach to filtering for potential anglicisms produces many false positives. Many German stop words in NLTK (Bird et al., 2009) overlap with English words from SpaCy’s (Honnibal and Montani, 2017) English dictionary and can be removed. We then used a word-to-word translator (Choe et al., 2020) with a smaller vocabulary than SpaCy or NLTK, to narrow down our list of anglicisms, and measured the length of this list. In our eight policies, we found 26 anglicisms on average, which impairs informational fairness by discriminating against demographics, where anglicisms are not commonly used.
@@
Table 1. Overview of surface measures
@@
words / policy 4809.59
paragraphs / policy 149.63
words / paragraph 32.14
headings / policy 33.82 (6.15% have no heading)
heading types 2.85
words / heading 4.83
lists 91.59 % of policies
other text formatting strong: 71.52 % of policies italics: 15.37 % of policies
Table 1 provides statistics on the document structure of all 618 policies. The average policy is very long. A dyslexic, who needs $\approx$2 minutes for 250 words (Musch and Rösler, 2011; Martelli et al., 2014), would need to spend almost 165 minutes reading the longest policy. The average reading time per policy for an average reader (Musch and Rösler, 2011) would be just under 20 minutes.
@@
4.3. Discussion
Text statistics, FRE and, to some extent, formatting have already been measured for privacy policies aiming at transparency (Bartelt and Buchmann, 2024). However, existing work did not use this information to assess key aspects of fairness. While informational fairness has previously been assessed for the predictions of machine learning models (Schoeffer et al., 2022), privacy policies are a different problem. Privacy policies are intended to balance the information asymmetry between provider and user (completeness), and add the aspect of potential discrimination due to a lack of inclusiveness in the presentation of information. Most approaches to privacy policies from the related work (cf. Sec. 2.3) only focus on completeness of information. Our preliminary results address different issues and allow for no comparison.
5. Representational Fairness
Our approach regarding representational fairness uses the methods outlined in Section 2.2.2 to assess the ability-related, socioeconomic, age-related, political, nationality-related, gender-related, sexuality-related, and cultural social / demographic axes.
5.1. Approach
Assessing representational fairness requires descriptor terms that capture various demographics (cf. Sec. 2). Based on the descriptor terms, a quantitative and semantic assessment can be carried out.
Fairness descriptor terms: We propose to carefully select existing descriptor terms, e.g., for gender (Muller et al., 2023), or via translation service from other languages (Smith et al., 2022). Finally, we propose to use a multilingual LLM to filter the resulting set, generate contextualized translations, or automatically generate policy-specific descriptor terms.
Quantitative and semantic fairness assessment: The descriptor terms enable us to estimate how often different groups are represented in a policy. Semantically, we propose to use the descriptor terms to assess the invariance of sentiment toward changed protected group membership by replacing existing entities with changed ones of varying protected groups. Finally, we propose to leverage the SAME metric (Schröder et al., 2021) for bias detection. We want to adapt approaches for toxicity detection (Risch et al., 2021), and debiasing the detection itself (Garg et al., 2023).
5.2. Preliminary Results
Again, our test case is based on German language privacy policies from the Top-100 German web shops. With this series of experiments, we evaluated the entire data set of 618 policies. We translated the descriptor terms from HolisticBias (Smith et al., 2022) to German with deep-translator (Baccouri, 2020) utilizing the Google Translator API, and we manually removed homonyms and other artifacts from the translation. We appended terms from the categories “nationality” and “political” that are specific for a German context, and we integrated German gender descriptor terms (Muller et al., 2023). We perform a quantitative assessment.
@@ 5.2. Preliminary Results
5.3. Discussion
To the best of our knowledge, representational fairness has not yet been assessed for privacy policies. We have only investigated quantitatively as a first step. For more nuanced and accurate findings, we want to assess semantics as a next step. This addresses false positives due to homonyms and has been used effectively in recent research (cf. Sec.  2.2.2).
6. Ethics and Morality
This assessment uses approaches from Section 2.2.3 to screen the processes, rights, and obligations declared in a privacy policy.
6.1. Approach
@@
Vulnerabilities and Proportionality: We propose to train a specific vulnerability classifier, similar to (Sánchez et al., 2021). Based on that, the proportionality of a policy depends on the business activities of the data holder. The assessment of proportionality requires weighing the difficult-to-quantify reasons for a case-specific process against its induced vulnerabilities. This raises challenges for typical ML classification models, so we suspect an ethicist’s judgment to be superior (cf. Section 2). That’s why we propose utilizing an LLM as a new approach to be tested. We use the classification result of the vulnerability classifier to prompt the LLM.
@@
General Ethics: We also suggest testing general ethics issues with an LLM, using a broad framing of the prompt. Because this is an explorative approach, we propose to request the LLM to state the criteria of its assessment first, and to quantify them on a five-point Likert scale in a subsequent step. We know from prompt engineering (Nori et al., 2023) that this can be achieved by using the entire policy as context and a task description in the prompt, and by asking the LLM to reflect its result. As a supplement, we suggest approaches for sentiment analysis (Guhr et al., 2020) on the LLM outputs, to get an approximate overall proxy for the ethics of a policy.
6.2. Preliminary Results
@@
Again, we use the German language policies from the German Top-100 web shops. For the time being, we focus on general morality and leave aside the other aspects. We created a prompt with a broad scope (see Figure 1), and tested it with GPT4turbo (OpenAI, 2023) on eight policies, one for each year in the data set. We asked the LLM not only to make a qualitative assessment, but also to rate it on a five-point Likert scale. Due to the lack of space, we only report the Likert ratings (Table 2).
@@
As the first column of Table 2 shows, we can aggregate the LLM assessments into 12 distinct criteria for the ethics of privacy policies. As desired, the criteria capture general ethical requirements on a broad scope and go beyond the GDPR, e.g., regarding the protection of minors. An LLM answers each prompt with a new internal seed, comparable to an expert asked to make an assessment from a different point of view. For this reason, we let the LLM evaluate each prompt in five runs. The second and third columns of Table 2 show in how many runs and policies the criteria appear. Some criteria just appeared in one of the runs for a policy. While integrating a fixed set of criteria into the prompt would enable a standardized assessment, running the LLM multiple times results in different aspects being considered, which has some merit for exploration. The fourth column shows the Likert-scores between 1 (worst) and 5 (best). Our most alarming findings are the usage of data, the (international) transmission, and sharing of data, as well as communication and changes.
System prompt: You are a critically scrutinizing, experienced expert in an ethics council on issues around fairness and justice and are committed to consumer protection for fairness. You have 20 years of experience in ethical issues around consumer protection, data protection, cybersecurity and related fairness aspects. You have already provided precise expert opinions in many court proceedings. Your task is to uncover aspects in privacy policies that are ethically questionable. You must clearly explain your thoughts and how you came to your conclusion. Explain step by step! Also disclose the ethical and moral principles on which you base your judgment. In particular, point out potential for improvement or risks to users. IMPORTANT: Also rate the privacy policy for each of your criteria on a 5-point Likert scale. Check whether the following privacy policy is fair and ethical towards its users. The user wants a critical review and an assessment of vulnerabilities in privacy policies. User prompt: The privacy policy: [Privacy Policy]
Figure 1. Prompt template (Square brackets replaced with privacy policy; German template translated to English)
Table 2. Overview of LLM ethics assessments
Criteria # Runs # Policies Average Score
transparency (precise, complete, comprehensible) 38 8 3.20
data subject control and autonomy (user rights, freedom of choice) 39 8 3.35
data minimalization and purpose binding 26 8 3.06
data usage (surveillance concerns, automated decisions) 21 8 2.71
data storage and deletion 8 4 3.38
data protection and security (avoiding misuse or leakage) 30 8 3.38
(international) data transmission and sharing (third-parties, where is what transferred) 27 8 2.74
compliance with data protection regulation (GDPR, relevant countries data protection regulation and standards) 9 7 3.56
communication and changes (timeliness, changes, risks) 8 5 2.63
protection of minors 4 2 3.50
fairness (no discrimination, fair use) 7 4 3.14
assurance, accountability and governance (compliance, reliability, assurance, ability to react) 9 6 3.22
6.3. Discussion
@@
Existing work has not automated ethics assessment so far, i.e., an LLM for ethics evaluation of privacy policies is a novel approach. Our findings are in line with related work (cf. Sec.  2.3) regarding problematic data use. We found the GDPR principles, which are hard ethics, represented in the ethics criteria, which we identified automatically with GPT4turbo. As we exemplarily investigated eight policies with a multilingual LLM, our list of criteria is just a starting point. There may be further criteria worth considering by utilizing a German LLM, more sophisticated prompting and a large and diverse set of policies.
7. Use Cases and Applications
@@
Our approach enables researchers to automatically scan a large corpus of privacy policies for such issues. This allows to obtain an overview of how widespread which kind of fairness issue is in privacy policies. Due to the ties between fairness and the GDPR, this might also help authorities to single out privacy policies that are problematic from a legal point of view. Our data set allows to analyze how such issues have developed or subsided over time. This provides an opportunity for researchers to investigate whether changes in legislation or major privacy events have led to fairer or less fair policies. Writing a fair policy is challenging. In the case of AI applications, for example, it is difficult even for experts to assess the personal risks associated with the use of personal data. Our approach allows to automatically identify aspects, that are associated with various kinds of bias, discrimination, etc. It might be useful to integrate this either into existing generators for privacy policies or to develop a stand-alone tool, to support data protection officers in writing complex, but fair privacy policies. Also, our work can be used as a foundation for data assessment for legal NLP models trained on privacy policies.
8. Conclusion